# Improving Adversarial Energy-Based Model via Diffusion Process

### v1

@@ 1 Introduction
Energy-based models (EBMs) are known as one type of generative models that draw inspiration from physics and have been widely studied in machine learning (Hopfield, 1982; Hinton & Sejnowski, 1983; Smolensky, 1986). EBMs define an unnormalized probability distribution over data space from a Gibbs density, which can be useful for several visual tasks, such as image classification (Grathwohl et al., 2020), out-of-distribution (OOD) detection (Liu et al., 2020) or semi-supervised learning (Gao et al., 2020). However, training an EBM using maximum likelihood estimation can be challenging due to the lack of a closed-form expression for the normalization constant. MCMC-based EBMs (Du & Mordatch, 2019; Nijkamp et al., 2019) evaluate the gradient of the objective through Markov chain Monte Carlo (MCMC) sampling on the defined energy function, which can be computationally expensive for both training and sampling. Adversarial EBMs (Grathwohl et al., 2021; Geng et al., 2021) introduce a ’generator’ to form a minimax game between alternative optimization of this generator and energy function, allowing for MCMC-free EBM training and fast sampling.
@@ 2 Related Work
There is a long history of EBMs in machine learning dating back to Hopfield networks (Hopfield, 1982) and Boltzmann machines (Hinton & Sejnowski, 1983). Recently, deep EBMs have grown in popularity, especially for image generation, and can broadly be classified into two categories. MCMC-based methods simulate Markov chains during training (Du & Mordatch, 2019; Arbel et al., 2021) or sampling (Song et al., 2020; Song & Ermon, 2019). This can be expensive, slow, and hard to control. Cooperative learning methods (Xie et al., 2020) jointly train a top-down generator and an energy function for MCMC teaching, using the generator as a fast initializer for Langevin sampling to alleviate the MCMC burden. They, however, remain inefficient.
@@
The original diffusion-based model (Ho et al., 2020; Sohl-Dickstein et al., 2015) learns a finite-time reversed diffusion from a forward process and gets strong performance on image generation, but the sampling process is slow. Several works (Song et al., 2021a; Lu et al., 2022b; Bao et al., 2022) improve training-free samplers and significantly improve the sampling speed going from 1000 steps sampling to only a few. DDGAN (Xiao et al., 2022) proposed to model each denoising step using a multimodal conditional GAN, tackling the generative learning trilemma of fast sampling while maintaining strong mode coverage and sample quality. However, it fails to provide a density estimate, which is crucial for many tasks.
Although some diffusion-based methods (Kingma et al., 2021; Song et al., 2021b; Lu et al., 2022a) use score functions or variational lower bounds to estimate the density of the data distribution, these methods tend to be complicated and inefficient, which is precisely where energy-based models excel. DRL (Gao et al., 2021) tried to combine EBM with a diffusion-based model and got state-of-the-art generation performance on several image datasets, but it relies on MCMC sampling which remains inefficient.
@@ 4 Experiments
We evaluate our DDAEBM in different scenarios across different data scales from 2-dimension toy datasets to large-scale image datasets. We test our energy function mainly on toy datasets and MNIST datasets which are easy to visualize and intuitive to measure. For large-scale datasets, we focus on image generation. We further perform some additional studies such as out-of-distribution (OOD) detection and ablation studies to verify our model’s superiority. We briefly introduce the network architecture design, while additional implementation details are presented in the Appendix. We will release the code alongside the paper. For the generator trained on large image datasets, we adopt the same modified NCSN++ architecture as DDGAN (Xiao et al., 2022) or DDPM++ with a slightly different structure as described in Score SDE (Song et al., 2021c), where diffused samples $\boldsymbol{x}_{t}$, time $t$ and latent variable $\mathbf{z}$ are the input of the network. For the energy function, we adopt traditional NCSN++ or DDPM++ architectures from Score SDE except that we remove the last scale-by-sigma operation and replace it with a negative $\ell_{2}$ norm between input $\boldsymbol{x}_{t}$ and the output of the Unet (Du et al., 2023), i.e.
@@ 4.2 Fitting flow models with energy function
4.3 Image generation
Refer to caption Refer to caption Refer to caption
To confirm the ability of our model to effectively scale to large-scale datasets, we conduct experiments of generation task on 32$\times$32 CIFAR-10 (Krizhevsky et al., 2009), 64$\times$64 CelebA (Liu et al., 2015), and 128$\times$128 LSUN church (Yu et al., 2015) datasets. Each dataset represents a significant increase in complexity. All datasets are scaled to $[-1,1]$ for pre-processing. For quantitative results, we adopt commonly used Fréchet inception distance (FID) and Inception Score (IS) to evaluate sample fidelity and the number of function evaluations (NFE) to evaluate sampling time. Table 2 shows quantitative results on CIFAR-10. We observe our model achieves FID 4.82 and IS 8.86, outperforming existing EBMs by a significant margin and performing comparably to strong baselines from other generative models such as GANs and diffusion-based models, which struggle with efficient density estimation.
@@ 4.3 Image generation
Note that our DDAEBM is still less performant than DDGAN even though they use the same generator structure and parameterization of the generated denoising model. This outcome is anticipated because the energy function in DDAEBM is tasked with providing a density estimate, not merely functioning as a discriminator. Therefore, the energy function in DDAEBM carries a heavier workload than the discriminator in DDGAN. In contrast, the discriminator in DDGAN merely distinguishes between real and fake sample pairs, primarily concentrating on the quality of the samples rather than on optimizing data likelihood.
@@
For CelebA, we only report FID scores in Table 3 since the IS score is not widely reported. Our model get the best performance among adversarial EBMs and outperforms score matching-based model NCSNv2 and VAE-based model NVAE. For LSUN church, Fig. 2 depicts high-fidelity synthesis sampled from the generator. We calculate FID on 50,000 samples using a PyTorch implementation from DDGAN and our model achieves 13.80 for this metric.
@@
DDAEBM 10.29
@@ 4.4 Out of distribution detection
The energy function of EBMs can be viewed as an unnormalized density function that assigns low values on out-of-distribution (OOD) regions and high values on data regions, which is suitable for detecting OOD samples. To test this, we first train our DDAEBM on CIFAR-10 and calculate the energy $E_{\theta}(\mathbf{x}_{0},0)$ on in-distribution images from the CIFAR-10 test set and out-of-distribution images from datasets including SVHN (Netzer et al., 2011), Texture (Cimpoi et al., 2014), CIFAR-100 (Krizhevsky et al., 2009) and CelebA. Following (Xiao et al., 2021), we use the area under the ROC curve (AUROC) as a quantitative metric on the energy scores, where high AUROC indicates that the model correctly assigns low density to OOD samples. Results are shown in Table 4, from where we can see our model achieves comparable performance with most of the baselines chosen from recent EBMs, VAEs, and Glow. Our model performs the best on SVHN and Texture datasets while on CIFAR-100 and CelebA it performs at a moderate level. Note that VAEBM performs slightly better than ours on most datasets, but it requires MCMC sampling to refine its generation, which can be inefficient. As is pointed out by Zhang et al. (2021), no method can guarantee better than random chance performance without assumptions on which out-distributions are relevant. Hence we see no clear winner on this task across all the datasets and we should take the results with a grain of salt.
@@
Fig 4 further show histograms of the energy output at $t=0$ between the CelebA test dataset and several OOD distributions. For fake data, we diffuse real data $\mathbf{x}_{0}$ for one step and use $p_{\phi}(\mathbf{x}_{0}|\mathbf{x}_{1})$ to generate fake data. For noisy data, we add noise to the real data with standard deviations of 0.01, 0.1, and 0.5. From Fig. 4 we observe our energy function assigns high values to real data and low values to fake and noisy data. Additionally, the energy values gradually decrease as the added noise increases. Although fake data and noisy data with 0.01 standard deviation have imperceptible differences for visualization, their energy values still exhibit clear distinctions when compared to real data. This demonstrates our energy function’s capability for OOD detection.
@@ 4.5 Ablation studies
First, we investigate the effects of some proposed modifications in our model including the latent variable in our defined denoising distribution, introduced posterior $q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$ and Jeffrey divergence for the generator’s training. Table 5 reports FID and IS scores on CIFAR-10 dataset and AUROC for the OOD task with CIFAR-10/SVHN as the in-distribution/out-of-distribution datasets. We replace sampling from $p(\mathbf{z})$ in Eq. (9) with all zero-vectors to remove the effect of $\mathbf{z}$ and observe the FID and IS scores are significantly worse. We also train our model with the most commonly used KL divergence and get similar performance on generation, but OOD performance significantly drops. We further remove $\log q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$-related term in our objective, which is equivalent to training a sequence of WGAN-0GPs embedded in a diffusion process. The generation and OOD performance are similar to the model trained with KL divergence. This experiment indicates the importance of our Jeffrey divergence and introduced $q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$ for the training of energy function. More results can be found in the Appendix.
@@ 5 Conclusion
Impact Statement.
@@ B.2 Network structure
Energy $permutate(\mathbf{x}_{t})$ $SE(t)$ 4 NICE layers NICE scale layer NICE layer FC 1000 PReLU ×$\times 4$ + FC 1000 PReLU ($SE(t)$) FC 392
@@ B.3 Hyperparameter settings
$\beta_{min},\beta_{max}$ in Eq. 47 (0.1, 10) (0.1, 20) (0.1, 20) (0.1, 20)
$w$, $w_{\text{mid }}$ in Eq. 21 (1, 1) (1, 1) (1, 1) (0.6, 0.2)
@@ C.1 More comparisons on Toy dataset
We also show more generation results in Fig 7 to compare our model with DDGAN and DDAEBM trained with KL divergence. Although DDGAN can achieve impressive sample quality on the 25-Gaussians dataset, as demonstrated in its initial paper, it gets limited quality on the Swiss roll dataset, which means DDGAN is not stable and robust enough across various datasets. Our DDAEBM can get much better generation performance on Swiss roll dataset, verifying our model’s advantage in terms of robustness. We can also observe that without the reversed KL divergence, DDAEBM fails to converge to each mode, leading to poor sample quality. Our DDAEBM trained with symmetric Jeffrey divergence can improve generation a lot, resulting in satisfactory sample quality on each mode. This experiment demonstrates besides the training of energy function as shown in Table 5, training with Jeffrey divergence also has superiority to the generator’s training by adding an extra reversed KL divergence.
@@ C.2 Mode counting
Generative Adversarial Networks (GANs) are notorious for mode collapse which is a phenomenon in which the generator function maps all samples to a small subset of the observation space. It’s well known that EBM can alleviate mode collapse because of its entropy term. Therefore, we also evaluate the mode coverage of our model on the StackedMNIST dataset. StackedMNIST is a synthetic dataset that contains images generated by randomly choosing 3 MNIST images and stacking them along the RGB channels. Hence the true total number of modes is 1,000, and they are counted using a pretrained MNIST classifier. Similar to DDGAN, we also follow the setting of (Lin et al., 2020) and report the number of covered modes and the KL divergence from the categorical distribution over 1000 categories of generated samples to true data in Table 11. It shows that our model covers all modes and achieves the lowest KL compared to GAN-based models and EBMs. This demonstrates our model is effective in mitigating mode collapse.

### v2

@@ 1 Introduction
Energy-based models (EBMs) are known as one type of generative models that draw inspiration from physics and have been widely studied in machine learning (Hopfield, 1982; Hinton & Sejnowski, 1983; Smolensky, 1986). EBMs define an unnormalized probability distribution over data space from a Gibbs density, which can be useful for several visual tasks, such as image classification (Grathwohl et al., 2020), out-of-distribution (OOD) detection (Liu et al., 2020) or semi-supervised learning (Gao et al., 2020). However, training an EBM using maximum likelihood estimation can be challenging due to the lack of a closed-form expression for the normalization constant. MCMC-based EBMs (Du & Mordatch, 2019; Nijkamp et al., 2019) evaluate the gradient of the objective through Markov chain Monte Carlo (MCMC) sampling on the defined energy function, which can be computationally expensive for both training and sampling. Adversarial EBMs (Grathwohl et al., 2021; Geng et al., 2021) introduce a generator to form a minimax game between alternative optimization of this generator and energy function, allowing for MCMC-free EBM training and fast sampling.
@@ 2 Related Work
There is a long history of EBMs in machine learning dating back to Hopfield networks (Hopfield, 1982) and Boltzmann machines (Hinton & Sejnowski, 1983). Recently, deep EBMs have grown in popularity, especially for image generation, and can broadly be classified into two categories. MCMC-based methods simulate Markov chains during training (Du & Mordatch, 2019; Pang et al., 2020; Arbel et al., 2021) or sampling (Song et al., 2020; Song & Ermon, 2019). This can be expensive, slow, and hard to control. Cooperative learning methods (Xie et al., 2020, 2021, 2022; Cui & Han, 2023) jointly train a generator and an energy function for MCMC teaching, using the generator as a fast initializer for Langevin sampling to alleviate the MCMC burden. They, however, remain inefficient.
@@
The original diffusion-based model (Ho et al., 2020; Sohl-Dickstein et al., 2015) learns a finite-time reversed diffusion from a forward process and gets strong performance on image generation, but the sampling process is slow. Several works (Song et al., 2021a; Lu et al., 2022b; Bao et al., 2022) improve training-free samplers and significantly improve the sampling speed going from 1000 steps sampling to only a few. DDGAN (Xiao et al., 2022) proposed to model each denoising step using a multimodal conditional GAN, tackling the generative learning trilemma of fast sampling while maintaining strong mode coverage and sample quality. However, its discriminator fails to provide a density estimate, which is crucial for many tasks (Du et al., 2023; Zhang et al., 2022; Du et al., 2022).
Although some diffusion-based methods (Kingma et al., 2021; Song et al., 2021b; Lu et al., 2022a) use score functions or variational lower bounds to estimate the density of the data distribution, these methods tend to be complicated and inefficient, which is precisely where energy-based models excel. DRL (Gao et al., 2021) and CDRL (Zhu et al., 2024) tried to combine EBM with a diffusion-based model and got state-of-the-art generation performance on several image datasets. Still, they relied on MCMC sampling which remains inefficient.
@@ 4 Experiments
We evaluate our DDAEBM in different scenarios across different data scales from 2-dimension toy datasets to large-scale image datasets. We test our energy function mainly on toy datasets and MNIST datasets which are easy to visualize and intuitive to measure. For large-scale datasets, we focus on image generation. We further perform some additional studies such as out-of-distribution (OOD) detection and ablation studies to verify our model’s superiority. We briefly introduce the network architecture design, while additional implementation details are presented in the Appendix. For the generator trained on large image datasets, we adopt the same modified NCSN++ architecture as DDGAN (Xiao et al., 2022) or DDPM++ with a slightly different structure as described in Score SDE (Song et al., 2021c), where diffused samples $\boldsymbol{x}_{t}$, time $t$ and latent variable $\mathbf{z}$ are the input of the network. For the energy function, we adopt traditional NCSN++ or DDPM++ architectures from Score SDE except that we remove the last scale-by-sigma operation and replace it with a negative $\ell_{2}$ norm between input $\boldsymbol{x}_{t}$ and the output of the Unet (Du et al., 2023), i.e.
@@ 4.2 Fitting flow models with energy function
4.3 Image generation
Refer to caption Refer to caption Refer to caption
To confirm the ability of our model to scale to large-scale datasets effectively, we conduct experiments of generation task on 32$\times$32 CIFAR-10 (Krizhevsky et al., 2009), 64$\times$64 CelebA (Liu et al., 2015), and 128$\times$128 LSUN church (Yu et al., 2015) datasets. Each dataset represents a significant increase in complexity. All datasets are scaled to $[-1,1]$ for pre-processing. For quantitative results, we adopt commonly used Fréchet inception distance (FID) and Inception Score (IS) to evaluate sample fidelity and the number of function evaluations (NFE) to evaluate sampling time. Table 2 shows quantitative results on CIFAR-10. We observe our model achieves FID 4.82 and IS 8.86, outperforming existing adversarial EBMs by a significant margin and performing comparably to strong baselines from other generative models such as GANs and diffusion-based models, which struggle with efficient density estimation.
@@ 4.3 Image generation
Note that our DDAEBM is still less performant than DDGAN even though they use the same generator structure and parameterization of the generated denoising model. We further add an additional denoising step (Song & Ermon, 2020) using our energy function to refine the generated samples, we find the FID score can be improved to 3.73, which is on par with DDGAN. This denoising step can be approximately viewed as a one-step MCMC using our energy function. With this minor modification, we successfully bridge the gap between DDGAN and DDAEBM in terms of generation performance. Besides, the energy function in DDAEBM is tasked with providing a density estimate, not merely functioning as a discriminator. Therefore, the energy function in DDAEBM carries a heavier workload than the discriminator in DDGAN. In contrast, the discriminator in DDGAN merely distinguishes between real and fake sample pairs, primarily concentrating on the quality of the samples rather than on optimizing data likelihood. This explains why the network of our energy function is stronger than the discriminator of DDGAN.
For CelebA, we only report FID scores in Table 3 since the IS score is not widely reported. Our model gets the best performance among adversarial EBMs and outperforms score matching-based model NCSNv2 and VAE-based model NVAE.
@@
DDAEBM (ours) 10.29
For LSUN church, Fig. 3 depicts high-fidelity synthesis sampled from the generator. We calculate FID on 50,000 samples using a PyTorch implementation from DDGAN and our model achieves 13.80 for this metric.
@@ 4.4 Out of distribution detection
The energy function of EBMs can be viewed as an unnormalized density function that assigns low values on out-of-distribution (OOD) regions and high values on data regions, which is suitable for detecting OOD samples. To test this, we first train our DDAEBM on CIFAR-10 and calculate the energy $E_{\theta}(\mathbf{x}_{0},0)$ on in-distribution images from the CIFAR-10 test set and out-of-distribution images from datasets including SVHN (Netzer et al., 2011), Texture (Cimpoi et al., 2014), CIFAR-100 (Krizhevsky et al., 2009) and CelebA. Following Xiao et al. (2021), we use the area under the ROC curve (AUROC) as a quantitative metric on the energy scores, where high AUROC indicates that the model correctly assigns low density to OOD samples. Results are shown in Table 4, from where we can see our model achieves comparable performance with most of the baselines chosen from recent EBMs, VAEs, and Glow. Our model performs the best on SVHN and Texture datasets while on CIFAR-100 and CelebA it performs at a moderate level. Note that VAEBM performs slightly better than ours on most datasets, but it requires MCMC sampling to refine its generation, which can be inefficient. As is pointed out by Zhang et al. (2021), no method can guarantee better than random chance performance without assumptions on which out-distributions are relevant. Hence we see no clear winner on this task across all the datasets and we should take the results with a grain of salt.
@@
Fig 4 further shows histograms of the energy output at $t=0$ between the CelebA test dataset and several OOD distributions. For fake data, we diffuse real data $\mathbf{x}_{0}$ for one step and use $p_{\phi}(\mathbf{x}_{0}|\mathbf{x}_{1})$ to generate fake data. For noisy data, we add noise to the real data with standard deviations of 0.01, 0.1, and 0.5. From Fig. 4 we observe our energy function assigns high values to real data and low values to fake and noisy data. Additionally, the energy values gradually decrease as the added noise increases. Although fake data and noisy data with 0.01 standard deviation have imperceptible differences for visualization, their energy values still exhibit clear distinctions when compared to real data. This demonstrates our energy function’s capability for OOD detection.
@@ 4.5 Ablation studies
First, we investigate the effects of some proposed modifications in our model including the latent variable in our defined denoising distribution, introduced posterior $q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$ and Jeffrey divergence for the generator’s training. Table 5 reports FID and IS scores on CIFAR-10 dataset and AUROC for the OOD task with CIFAR-10/SVHN as the in-distribution/out-of-distribution datasets. We replace sampling from $p(\mathbf{z})$ in Eq. (9) with all zero-vectors to remove the effect of $\mathbf{z}$ and observe the FID and IS scores are significantly worse. We also train our model with the most commonly used KL divergence and get similar performance on generation, but OOD performance significantly drops. We further remove $\log q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$-related term in our objective, which is equivalent to training a sequence of WGAN-0GPs embedded in a diffusion process. The generation and OOD performance are similar to the model trained with KL divergence. This experiment implies that asymmetric KL divergence and lack of $\log q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$ lead to inaccurate learning of the energy function which, in turn, affects e.g. OOD performance. Symmetric Jeffrey divergence and entropy term $q_{\psi}(\mathbf{z}|\mathbf{x}_{t-1},\mathbf{x}_{t})$ ensure a better fit between the generated distribution and the energy function. This improves the energy training, resulting in improved OOD detection. More results can be found in the Appendix.
@@ 5 Conclusion
Acknowledgements
This work was supported by a research grant (42062) from VILLUM FONDEN. This project received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement 757360). The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606).
Impact Statement
@@ B.2 Network structure
Energy
$permutate(\mathbf{x}_{t})$ $SE(t)$
4 NICE layers
NICE scale layer
NICE layer
FC 1000 PReLU ×$\times 4$
+ FC 1000 PReLU ($SE(t)$)
FC 392
@@ B.3 Hyperparameter settings
$\beta_{min},\beta_{max}$ in Eq. (47) (0.1, 10) (0.1, 20) (0.1, 20) (0.1, 20)
$w$, $w_{\text{mid }}$ in Eq. (21) (1, 1) (1, 1) (1, 1) (0.6, 0.2)
@@ C.1 More comparisons on Toy dataset
We also show more generation results in Fig. 7 to compare our model with DDGAN and DDAEBM trained with KL divergence. Although DDGAN can achieve impressive sample quality on the 25-Gaussians dataset, as demonstrated in its initial paper, it gets limited quality on the Swiss roll dataset, which means DDGAN is not stable and robust enough across various datasets. Our DDAEBM can get much better generation performance on Swiss roll dataset, verifying our model’s advantage in terms of robustness. We can also observe that without the reversed KL divergence, DDAEBM fails to converge to each mode, leading to poor sample quality. Our DDAEBM trained with symmetric Jeffrey divergence can improve generation a lot, resulting in satisfactory sample quality on each mode. This experiment demonstrates besides the training of energy function as shown in Table 5, training with Jeffrey divergence also has superiority to the generator’s training by adding an extra reversed KL divergence.
@@ C.2 Mode counting
Generative Adversarial Networks (GANs) are notorious for mode collapse which is a phenomenon in which the generator function maps all samples to a small subset of the observation space. It’s well known that EBM can alleviate mode collapse because of its entropy term. Therefore, we also evaluate the mode coverage of our model on the StackedMNIST dataset. StackedMNIST is a synthetic dataset that contains images generated by randomly choosing three MNIST images and stacking them along the RGB channels. Hence the true total number of modes is 1,000, and they are counted using a pretrained MNIST classifier. Similar to DDGAN, we also follow the setting of (Lin et al., 2020) and report the number of covered modes and the KL divergence from the categorical distribution over 1000 categories of generated samples to true data in Table 11. It shows that our model covers all modes and achieves the lowest KL compared to GAN-based models and EBMs. This demonstrates our model is effective in mitigating mode collapse.